# NewsInterview: a Dataset and a Playground to Evaluate LLMs’ Grounding Gap via Informational Interviews

### v1

@@ 4.2 Gameplay Design
To prepare our dataset for use in the simulated game environment, we group together: (1) source responses and ask an LLMLlama-3.1-70b to summarize a set of specific informational items and (2) interviewer questions and ask an LLM to summarize them into a set of high-level objectives. The sources’ informational items mimic the knowledge a source likely had going into the interviewManual evaluation confirms that these information items were present in the initial interview, are concise and non-overlapping. and the interviewer’s objectives represent the agendas they had prior to the conversationManual validation with professional journalists confirms that these outlines reasonably capture what a journalist might prepare before an interview and do not leak information.. Both of these summaries are represented in Figure 3 as Given, and are designed to give the interviewer-LLM and the source-LLM a basis for communication. For further examples of both, see Tables 7 and 6 in the Appendix.
Source Design: Personas and Persuasion
@@
Now, we introduce the design of the source. We focus attention on this construction to build a robust game environment that accurately mimics human interactions. To make gameplay varied and challenging, we draw from Sedorkin (2015) to design eight different personas: Anxious, Avoidant, Adversarial, Defensive, Straightforward, Poor Explainer, Dominating and Clueless. To see detailed descriptions of each persona, as well as example responses that the persona might give, see Table 3. Different personas give us the ability to study how interviewers perform in a wider array of scenarios and are designed to more carefully capture different challenges that journalists face.
Hardest Intermediate Easiest
@@
Table 2: Performance of LLMs as Interviewers, with Ablations Percentage of information items extracted (Reward %) in each interview by different language models (gpt-4o-mini, gpt-4o, Llama-3.1-70b, and Llama-3.1-8b) across three conditions: (1) Hardest: The full game, with information dependent on persuasion and persona. (2) Medium: an ablation removing the sources’ responsiveness to persuasion. (3) Easy: An ablation removing the random withholding of information (i.e. a source returns all relevant information items at each turn). We observe, perhaps unsurprisingly, that removing the source’s ability to withhold information (Intermediate $\rightarrow$ Easy) drastically increases the reward % at the end of the game. The removal of persuasion strategies has a smaller effect, with some models showing marginal gains (e.g., Llama-3.1-8b) and others slight losses (e.g., gpt-4o). This indicates that vanilla LLMs are poorly suited to this persuasion task.
@@
Figure 4: Comparison of gpt-4o’s performance across different persona types: extraction rewards and persuasion levels. The Adversarial type is by far the hardest to extract information from, however, it is easier to persuade. This could be because the LLMs are the most thrown off by adversarial sources.
[…]
@@
(b) Percentage (%) of Reward, by total reward, over time.
@@
Finally, based on the assessed persuasion level (1–5) of the conversation, we implement getItemsToReturn. This function takes in all relevant information items and randomly draws from a Beta distribution to determine what percentage of relevant information items to return. We choose 5 different parameterizations per persona, each corresponding to a different persuasion level. As can be seen in Figure 3, we choose these parameterizations such that the more persuaded a source is, the more left-skewed the distribution is. Each persona has a slightly different parameterization, reflecting that some personas need less persuasion (e.g. “Dominant”) while others do not drastically change how much information they return even with more persuasion (e.g. poor explainer). See Figure 6 in the Appendix for the Beta distributions for each source.
Source and Interviewer Responses
To implement the conversational components of our game, we stick with a minimalist setup, as LLMs have been well-observed to be proficient at basic dialogue given sufficient structure Ou et al. (2024). For Source, we provide the source with the information items to return, a description of its persona, and its prior assessed comfort/persuasion level. For Interviewer, we implement a prompting-based setup and leave more involved designs for future.
@@ 4.3 Game Simulation Results
We run our simulation for 200 interviews with four models as the interviewer: gpt-4o, gpt-4o-mini, Llama-3.1-70b and Llama-3.1-8b. For a source LLM, we used gpt-4o across all personas. Table 2 compares the performance of LLMs across three conditions: the full game, a version without persuasion, and a version where sources do not withhold information. In the full game, where sources’ responsiveness depends on persuasion and persona, the gpt-4o model performs the best, at 50.4%. However, when persuasion is removed, performance only marginally improves across all models (e.g., Llama-3.1-70b reaches 45.5%, while gpt-4o remains stable at 49.8%), indicating that other aspects of the game (i.e. inferring which information the source has witheld) also pose a challenge. In the easiest condition, where no information withholding occurs, all models perform significantly better, with reward percentages reaching over 80%, showing that withholding strategies are a major obstacle for current LLMs.
Figure 4(a) highlights the performance of gpt-4o across different source personas. The model achieves the highest information extraction from straightforward personas, while adversarial and defensive personas are the most challenging. Despite being harder to extract information from, adversarial sources are easier to persuade, as shown in Figure 4(b).
@@ 5 Discussion
Our results from Sections 3.3 and 4.3, taken together, show that news interview transcripts are a valuable and real-world dataset for learning about persuasion, grounding dialogue and multi-turn strategy. Our work builds off previous work identifying grounding gaps in language models Shaikh et al. (2024) and an emerging thread in gameplay Wongkamjan et al. (2024); Liu et al. (2024), by extending both directions into the news domain, where real-world data is plentiful.
In Section 3.3, we show that grounding persuasive dialogue exists in human interviewers, which LLMs fail to mimic. In Section 4.3 we show that LLMs struggle to extract information from sources embodying a variety of personas. Interestingly, particular personas, like adversarial or avoidant sources, are noticeably more difficult for the LLM extract information from. Figure 4(b), for instance, might suggest that the LLM models are more responsive to engagement tactics when sources display hostility rather than indifference or avoidance. In the context of prior work Liu et al. (2024); Chawla et al. (2021), which tends to examine all interactions from a singular persona, these distinctions can point to further investigations illuminating why certain personas are worse than others.
The ultimate solution for a propensity towards multi-turn strategies and grounding language in LLMs, we feel, is to incorporate longer-range reward signals into the training process Li et al. (2016). After all, grounding exists for a purpose: to influence an outcome borne out over many conversational turns Clark (1996); Cialdini (2009) (e.g. in therapeutic environments, the patient is more open and makes more progress Bohart (1999); in educational environments the student is more engaged and learns more Brown (1994)).
Our NewsInterview effort provides a game-playing framework to provide such a reward and test such strategies in a real-world scenario. This game-playing environment is complicated: throughout an interview, the LLM has to reason about what information it already has, what likely information the source has, and how to persuade the source to divulge this information. Each of these components is likely challenging for present-day LLMs. Figure 5(a) suggests that LLMs struggle to maintain effective information extraction as the conversation progresses, possibly due to difficulties in adjusting their questioning strategies over time. Furthermore, we see only marginal gains with large modeling sizes, suggesting that scale alone may not make meaningful progress.
Despite the considerable challenges this game imposes, NewsInterview is still simpler than other game-playing environments: agent-based strategy games typically require the agent to develop a strategy while other players are also developing strategies Chawla et al. (2021); Liu et al. (2024); Wongkamjan et al. (2024). In NewsInterview, the source is not typically trying to persuade the journalist, simply to decide how much information to divulge. These components, taken together, suggest that NewsInterview is an ideal real-world starting place to start to incorporate multi-turn strategic dialogue from LLMs. We imagine that agents built off of this training environment will one day be able to: (1) identify persona characteristics easily in other game-players (2) develop strategy books, similar to those humans use Sedorkin (2015); Harcup (2015) (3) reason about early-vs-late stage conversational tactics. These agents can not only help in practical tasks (e.g. by automating interviewing or by providing students an environment to learn in) but also give us more fundamental insights into the strategic approaches that work well.
Our work has to be viewed with the following fundamental limitations. getPersuasionLevel serves as the core component in our game-design that affects the strategy an agent can learn and deploy; thus, any flaw in this function can lead us to a suboptimal agentic strategy. Despite observing high correlations between humans and LLM assessments of persuasion, there could be many common scenarios we do not observe. For instance, there may be pathological degenerate cases: if, for instance, getPersuasionLevel returns a persuasion level of $5$ for agents that say “thank you”, then agents will learn to say this at every turn, and will not develop any other strategy. We implemented our personas using prompts outlining what different personas would say and what they would be persuaded by, as described in Section 4.2, under the expectation that the source-LLM would be able to generalize. Table 2 does indicate that the interviewer-agent did not simply overfit to certain words. However, if, in future work, we observe collapsing interviewer-agent behavior especially in the presence of trained approaches, we must be prepared to do more work to validate or improve getPersuasionLevel.
Additionally, future work could also focus on introducing additional reward signals into the game-play. Some pieces or information are more important than others. By analyzing how much time is spent discussing each of a source’s pieces of information, we might reward the agent more for the more important extractions. Another reward signal could be the quotability of the response. Often, an interviewer will seek to obtain specific, memorable lines to write about throughout the interview. There is line of research that tries to predict which parts of a text or speech are the most quotable Bendersky and Smith (2012); Lee et al. (2016); MacLaughlin et al. (2018). We believe both of these signals might be an interesting lens through which to improve NewsInterview.
@@ 6 Related Work
Recent research on large language models (LLMs) has highlighted a lack of grounding language and strategic dialogue Shaikh et al. (2024); Wongkamjan et al. (2024). Grounding communication, which involves the use of acknowledgments and affirmations to foster understanding and trust Clark (1996); cho-may-2020-grounding, is essential in various conversational settings such as education Kasneci et al. (2023), mental health Carlbring et al. (2023), and conflict resolution Argyle et al. (2023). Prior efforts to study and improve LLMs’ grounding dialogue capabilities have faced limitations due to the scarcity of large-scale, naturalistic datasets Kasneci et al. (2023). Existing datasets are either generated via crowdsourcing, which can result in unnatural dialogues Rashkin et al. (2019); Wang et al. (2019); Liu et al. (2021), or are small-scale due to privacy concerns, as in educational Caines et al. (2020) and therapeutic settings Gratch et al. (2014); Casey (2004).
Grounding dialogue exists to foster longer-term goals. LLMs have been shown to lack multi-turn strategic thinking and planning, which are critical for effective dialogue Chawla et al. (2023). While some studies have explored the use of game-playing environments to improve these aspects Li et al. (2016); Wongkamjan et al. (2024), and the development of ‘gamified’ persona-based subjects has a long history in dialogue agent research Colby (1981), there is still a need for realistic datasets and simulations that can facilitate the development of agents with longer-horizon rewards.
Research has increasingly focused on using game-playing environments to train agents in strategic dialogue. Lewis et al. (2017) introduced a negotiation task where agents learn to negotiate over resources through dialogue, demonstrating the potential for agents to develop complex negotiation strategies. Similarly, He et al. (2018) proposed a decoupled policy optimization approach for strategic dialogues in games. More recent works have extended strategic dialogue to complex, multi-agent environments. Gray et al. (2020) developed agents capable of playing the game Diplomacy at a human level by engaging in strategic planning and negotiation with other agents. Perolat et al. (2022) introduced a model for playing Stratego, showcasing multi-agent reinforcement learning in adversarial settings. Chawla et al. (2021) presented CASINO, a contextualized dialogue dataset designed for building negotiation agents, emphasizing the importance of context and strategy in dialogue systems. These works highlight the effectiveness of simulated game environments in training agents for strategic multi-turn dialogues, underscoring the potential of such approaches in enhancing LLMs’ strategic dialogue capabilities.
Despite these advancements, several gaps remain. Most existing models focus on task-specific dialogues with singular personas and lack the flexibility to handle a variety of conversational contexts and topics. Another gap is the scarcity of large-scale, naturalistic datasets to create a broader variety of settings for these games, hampering the development of more generalized and robust models Liu et al. (2021). Addressing these gaps is crucial for advancing the development of agents capable of engaging in sophisticated strategic dialogues akin to human interactions.
Our work addresses these gaps by introducing a large-scale dataset of 40,000 two-person informational interviews from NPR and CNN, providing a rich resource for studying grounding communication in a naturalistic setting. Additionally, we develop a realistic simulated environment, NewsInterview, which incorporates source personas and persuasive elements to evaluate and improve LLMs’ strategic dialogue capabilities.
@@ 7 Conclusion
In this paper, we have introduced a high-quality dataset of 40,000 two-person informational interviews from NPR and CNN, addressing the scarcity of large-scale dialogue data necessary for studying grounding communication. Our detailed discourse analysis reveals significant differences between LLM-generated dialogues and human interviewers, particularly in the use of grounding language and question types. Motivated by observation that long-term objectives guide turn-level grounding, we develop a realistic game environment, NewsInterview, to test and improve dialogue agents in informational interviews. Our experiments demonstrate that while source LLMs can mimic human behavior in information sharing, interviewer LLMs struggle with recognizing when questions are answered and engaging persuasively, leading to suboptimal information extraction. These findings underscore the need for enhancing LLMs’ strategic dialogue capabilities, and we believe that our dataset and simulation environment provide valuable resources for future research in this area.
@@ 8.1 Privacy and Ethical Considerations
All data used in this study are publicly available and do not contain personally identifiable information beyond what has been already made public by the news organizations. We adhere to ethical guidelines for data use and ensure that our processing respects the rights and privacy of individuals involved as well as the news organizations that collected this data. Since the dataset we create is derived from interviews that have already been published in academic settings, we believe we are infringing upon the copyright of the news organizations this data originally belonged it. Aside from ownership questions, there are still inherent risks in the use of real-world interview data. Some interviews might involve sensitive topics, and the ethical implications of using such data for model evaluation warrant careful consideration.

### v2

@@ 4.2 Gameplay Design
To prepare our dataset for use in the simulated game environment, we group together: (1) source responses and ask an LLM. [^8] Llama-3.1-70b to summarize a set of specific informational items and (2) interviewer questions and ask an LLM to summarize them into a set of high-level objectives. The sources’ informational items mimic the knowledge a source likely had going into the interview [^9] Manual evaluation confirms these information items are present in initial interviews and are non-overlapping. and the interviewer’s objectives represent the agendas they had prior to the conversation. [^10] Manual validation with professional journalists confirms that these outlines reasonably capture what a journalist might prepare before an interview and do not leak information. Both of these summaries are represented in Figure 3 as Given, and are designed to give the interviewer-LLM and the source-LLM a basis for communication. For further examples of both, see Tables 9 and 10in the Appendix.
Source Design Element #1: Personas
@@
Now, we introduce the design of the source. We focus attention on this construction to build a robust game environment that accurately mimics human interactions. To make gameplay varied and challenging, we draw from Sedorkin (2015) to design eight different personas: Anxious, Avoidant, Adversarial, Defensive, Straightforward, Poor Explainer, Dominating and Clueless. For descriptions of each persona, as well as example responses, see Table 11. These personas allow us to study how interviewers perform in a wider array of challenging scenarios.
Source Design Element #2: Persuasion
The following three functions, in sequence, power our gameplay: $\texttt{getRelevantInfoItems}\rightarrow\texttt{getPersuasionLevel}\rightarrow\texttt{getItemsToReturn}$. The first, getRelevantInfoItems, takes the interviewer’s question and determines which of the sources’ information items are most relevant; it is simply a retrieval function that we implement using an LLM. getPersuasionLevel is a function that determines the selected source’s level of comfort or persuasion (on a five point scale) in the current conversation. getItemsToReturn is a stochastic engine: it randomly selects, based on the persuasion level, the number of relevant information items to return: the more persuaded a source is, the more likely they are to return more information. The persuadability component to our gameplay increases the multi-turn strategy: because persuasion is assessed with reference to the entire interview, the interviewer gets more reward for spending words early in the interview persuading the source to feel comfortable.
Is it sound for the source-LLM to assess its own level of persuasion? As recent research has found, LLMs are poor detectors of when they are being persuaded Sakurai and Miyao (2024) and can even unknowingly persuade themselves Zeng et al. (2024). Furthermore, persuadability varies from person to person Wang et al. (2019); Hirsh et al. (2012). Luckily, source-persuasion is a well-studied field in journalism. As a starting point, we draw from Sedorkin (2015), and carefully design prompts asking an LLM to rate the persuasiveness of a prior conversation. Different source personas, according to Sedorkin (2015), are persuaded by different communication patterns: e.g., Anxious sources are distrustful of journalists; they are usually persuaded by phrases like “I will be as fair as possible.” We validate this in Section 4.3.
Source and Interviewer Responses
Based on the assessed persuasion level (1–5) of the conversation, we implement getItemsToReturn. This function takes in all relevant information items and randomly draws from a Beta distribution to determine what percentage of relevant information items to return. We choose five different parameterizations per persona, each corresponding to a different persuasion level. As can be seen in Figure 3, we choose these parameterizations such that the more persuaded a source is, the more left-skewed the distribution is. Each persona has a slightly different parameterization, reflecting that some personas need less persuasion (e.g., “Dominant”) while others do not drastically change how much information they return even with more persuasion (e.g., poor explainer). See Figures 6 and 7 in App. for the Beta distributions for each source.
4.3 Gameplay Validation
We conducted human trials to validate how well our gameplay environment approximates real interviews, focusing on persuasion as a pivotal dimension. Five participants, including two professional journalists and one journalism student, each served as the “source,” rating their own persuasion levels turn-by-turn on a five-point scale across 72 trials (576 turns total). The game’s LLM-based source also generated persuasion estimates. We found a moderate but significant correlation of $r=0.43$ (p < .0001). Excluding adversarial personas, correlation rose to $r=0.68$. Bootstrapped estimates confirmed the consistency of these results, and a power analysis following guidelines from Card et al. (2020) showed our sample size was adequate to detect this effect.
These trials center on persuasion because the other components of our source design (i.e., retrieval of correct informational items), while crucial, leverage prior, well-studied phenomena in retrieval-augmented LLMs and prompt engineering Lewis et al. (2020); OpenAI . Our environment reuses standard cross-encoder reranking and chain-of-thought prompts Wei et al. (2022); Reimers and Gurevych (2019), meaning that the correct factual content is generally well-handled without substantial new techniques. Minimal forms of self-reflection Shinn et al. (2023) were used to mitigate hallucinations, and no significant factual drift was observed. Hallucinations are well-studied in the literature Huang et al. (2025).
Taken together, this validation suggests that modeling source persuadability in a turn-level simulation is reasonably accurate and stable. By capturing how LLMs adapt their strategies across different personas and persuasion thresholds, our system can potentially serve as a stepping stone for training more sophisticated interview agents or supporting journalism students. Future work might expand the environment’s human trials, repeat experiments at larger scale, and incorporate further realism checks to ensure robust dialogue performance and fidelity.
Hardest Medium Easiest
@@ 4.3 Gameplay Validation
Table 2: Performance of LLMs as Interviewers, with Ablations Percentage of information items extracted (Reward percentage) in each interview by different language models (gpt-4o-mini, gpt-4o, Llama-3.1-70b, and Llama-3.1-8b) across three conditions: (1) Hardest: The full game, with information dependent on persuasion and persona. (2) Medium: an ablation removing the sources’ responsiveness to persuasion. (3) Easy: An ablation removing the random withholding of information (i.e., a source returns all relevant information items at each turn). We observe, perhaps unsurprisingly, that removing the source’s ability to withhold information (Medium $\rightarrow$ Easy) drastically increases the reward percentage at the end of the game. The removal of persuasion strategies has a smaller effect, with some models showing marginal gains (e.g., Llama-3.1-8b) and others slight losses (e.g., gpt-4o). This indicates that vanilla LLMs are poorly suited to this persuasion task.
@@
Figure 4: Comparison of gpt-4o’s performance across different persona types. The Adversarial type is by far the hardest to extract information from, however, it is easier to persuade. LLMs might be most the thrown off by adversarial sources.
@@
(b) Percentage (%) of Reward, by total reward.
@@ 4.4 Game Simulation Results
We run our simulation for 450 interviews with four LLMs as the interviewer111111gpt-4o, gpt-4o-mini, Llama-3.1-70b and Llama-3.1-8b and gpt-4o for the source-LLM across all personas. Table 2 compares the performance of LLMs across three conditions: the full game, a version without persuasion, and a version where sources do not withhold information. In the full game, where sources’ responsiveness depends on persuasion and persona, the gpt-4o model performs the best, at 50.4%. However, when persuasion is removed, performance only marginally improves across all models (e.g., Llama-3.1-70b reaches 45.5%, while gpt-4o remains stable at 49.8%), indicating that other aspects of the game (i.e., inferring which information the source has witheld) also pose a challenge. In the easiest condition, where no information withholding occurs, all models perform significantly better, with reward percentages reaching over 80%, showing that withholding is a major obstacle.
Figure 4(a) highlights the performance of gpt-4o across different source personas. The model achieves the highest information extraction from straightforward personas, while adversarial and defensive personas are the most challenging. Despite being harder to extract information from, adversarial sources are easier to persuade (Figure 4(b)).
@@ 5 Discussion
Our findings indicate that news interview transcripts provide a powerful, real-world resource for studying persuasive, grounding, and multi-turn strategies in dialogue systems. In particular, we build on prior work that highlights grounding gaps in large language models (LLMs) Shaikh et al. (2024a), extending insights from gameplay-inspired multi-turn dialogue research Wongkamjan et al. (2024); Liu et al. (2024) into a domain abundant with authentic data. By examining human interviewers’ behaviors, we illustrate how grounding and persuasion manifest naturally in real-world news interviews, yet remain difficult for current LLMs.
We show in Section 3.3 that humans consistently employ grounding dialogue throughout their interviews, a tactic LLMs fail to emulate effectively. In Section 4.4, we demonstrate how LLMs struggle to extract information from diverse source personas, particularly when those personas exhibit adversarial or avoidant traits. These findings underscore the significance of persona mismatches: while existing game-based dialogue studies often assume a single persona per environment Chawla et al. (2021); Liu et al. (2024), our results suggest that personae with different levels of hostility or indifference pose unique challenges for current models.
One way to address these limitations is to incorporate long-range reward signals during model training Li et al. (2016). Grounding dialogue and persuasion are inherently long-horizon phenomena Clark (1996); Cialdini (2009). In contexts like therapy, for instance, effective grounding fosters patient openness and lasting progress Bohart and Tallman (1999); in education, it encourages students’ sustained engagement and deeper learning Brown and Campione (1994). Our NewsInterview framework addresses this by providing an environment in which LLMs must continually strategize about which questions to ask, what information gaps need filling, and how to persuade sources to disclose details. This game-playing setting is still less complex than fully adversarial multi-agent domains Chawla et al. (2021); Liu et al. (2024); Wongkamjan et al. (2024) because the source’s goal is not to mislead but to selectively withhold information. Yet, even in this simplified scenario, LLMs struggle to maintain effective information extraction over multiple turns, pointing to deeper issues in question-asking. Future directions include refining our getPersuasionLevel function, introducing importance-weighted or quote-centric reward signals, and further validation.
@@ 6 Related Work
Research on large language models (LLMs) has increasingly underscored the importance of grounding language and strategic dialogue Shaikh et al. (2024a); Wongkamjan et al. (2024), especially in educational, mental health, and conflict resolution settings Clark (1996); Cho and May (2020); Kasneci et al. (2023); Carlbring et al. (2023); Argyle et al. (2023). However, current datasets for studying these phenomena are either crowdsourced and thus somewhat contrived Rashkin et al. (2019); Wang et al. (2019); Liu et al. (2021), or too small to capture the complexity of real-world interactions due to privacy constraints Gratch et al. (2014); Casey (2004); Caines et al. (2020). This lack of large-scale, naturalistic data limits progress on LLMs that effectively employ affirmations and acknowledgments to establish rapport and longer-term strategic thinking Kasneci et al. (2023).
Meanwhile, recent research has explored the use of simulated or game-based environments for training dialogue agents on multi-turn strategic planning Lewis et al. (2017); He et al. (2018); Gray et al. (2020); Perolat et al. (2022); Chawla et al. (2021). Though these environments can foster negotiation and strategy, they often rely on narrowly-defined tasks or singular personas. Consequently, there is still a need for broader conversational contexts and larger, more naturalistic datasets Liu et al. (2021). We address these gaps by introducing 40,000 two-person informational interviews from NPR and CNN, contributing a new large-scale dataset for studying grounding communication and strategic dialogue in realistic settings. Additionally, our NewsInterview simulation environment incorporates source personas and persuasive tactics, further advancing the development and evaluation of LLMs in long-horizon, multi-turn dialogues.
@@ 7 Conclusion
In this paper, we have introduced a high-quality dataset of 40,000 two-person informational interviews from NPR and CNN, addressing the scarcity of large-scale dialogue data necessary for studying grounding communication. Our detailed discourse analysis reveals significant differences between LLM-generated dialogues and human interviewers, particularly in the use of grounding language and question types. Motivated by observation that long-term objectives guide turn-level grounding, we develop a realistic game environment, NewsInterview, to test and improve dialogue agents in informational interviews. Our experiments demonstrate that while source-LLMs can mimic human behavior in information sharing, interviewer-LLMs struggle with recognizing when questions are answered and engaging persuasively, leading to suboptimal information extraction. These findings underscore the need for enhancing LLMs’ strategic dialogue capabilities.
@@ 8.1 Privacy and Ethical Considerations
All data used in this study are publicly available and do not contain personally identifiable information beyond what has been already made public by the news organizations. We adhere to ethical guidelines for data use and ensure that our processing respects the rights and privacy of individuals involved as well as the news organizations that collected this data. Since the dataset we create is derived from interviews that have already been published in academic settings, we believe we are not infringing upon the copyright of the news organizations this data originally belonged it. Aside from ownership questions, there are still inherent risks in the use of real-world interview data. Some interviews might involve sensitive topics, and the ethical implications of using such data for model evaluation warrant careful consideration.